# QUILL: Quotation Generation Enhancement of Large Language Models

### v1

@@ 1 Introduction
Figure 1: An example of Quotation Generation (QG) Evaluation. In existing QG tasks, LLMs are often prone to quotation hallucination, inconsistent context, lack of paragraph coherence, and lack of novelty in quotations. LLMs currently face great challenges in QG tasks.
@@ 4 Evaluation System for QG
Quotation Authenticity. Authenticity of quotations is crucial as it ensures the reliability and credibility of information Kington et al. (2021). To verify the authenticity of the quoted celebrity quotes, our study first employs search engines (such as Google Scholar [^1] https://scholar.google.com/ and Baidu Scholar [^2] https://xueshu.baidu.com/) to retrieve quotes and gather relevant search results. It’s crucial to ensure the accuracy and relevance of keywords during retrieval, aiming for representative results. Next, ChatGPT OpenAI (2022) is used to analyze and extract information from the search results, identifying key details related to the quotes, such as celebrity names, quote content, and sources (as illustrated in Figure.1). Finally, based on the extracted information, we verify whether the quote genuinely originates from the specific celebrity. If multiple sources are found, a manual comparison of authoritative sources (such as academic papers, biographies, and reputable media reports) is necessary to confirm the authenticity of the citations. The final score for quotation authenticity is defined as follows:
@@
Quotation Credibility. Generally speaking, in the context of quoting, the source of the quote will be mentioned, such as the author, classic literature, or other sources. Ensuring consistency between the citation and the mentioned author or source is crucial for maintaining contextual coherence and information accuracy Aly et al. (2024). In order to confirm whether the citation meets the source restriction mentioned in the context, our study first extracts the source restriction of the context, and then compares and analyzes it with the extraction result of the previous indicator. If the source matches, the citation is marked as trustworthy, as shown in Fig.1. The final score for quotation credibility is defined as follows:
@@
Semantic Matching. Improper quotation may lead to misunderstandings or misinterpretations of the original meaning, thereby weakening the effectiveness and persuasiveness of the argument Quora (2020). To assess the consistency between citations and their contexts, our study employs two advanced language models, Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023). These models are used to calculate the PPL of the right context given the prior text and the citation quote, respectively.
Then, we take the average of the perplexity values calculated by the two models as the semantic matching score, which can balance the judgment of the two models and reduce the possible deviation caused by a single model. If the metric score is low, it indicates that the quoted quote is highly semantically consistent with the context, otherwise the rationality of the quote needs to be reconsidered.
In summary, the calculation formula for semantic matching is as follows:
$PPL_{m}=sum\left(PPL_{q}\left(c_{r}\mid[c_{l};q]\right)+PPL_{l}\left(c_{r}\mid[c_{l};q]\right)\right)$ (4)
where $\text{PPL}_{q}\left(c_{r}\mid[c_{l};q]\right)\text{, }\text{PPL}_{l}\left(c_{r}\mid[c_{l};q]\right)$ represent the perplexity of Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023) respectively.
However, for the convenience of calculation, we set a benchmark PPL value (PPL=50), which is usually the target value of model performance or a reasonable upper limit, and combined it with the feature that the lower the PPL, the higher the semantic matching degree. The final score for semantic fluency is as follows:
$S_{m}=\left(1-\frac{PPL_{m}}{50}\right)*100\%$ (5)
@@
Semantic Fluency. After quotation, it is necessary to ensure that the entire context is fluent and coherent to maintain semantic consistency and logical integrity. In order to measure the fluency of the entire context after quotation, we use the same models as the previous metric, Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023), to calculate the perplexity of the entire context. Similarly, we also take the average of the two models as the semantic fluency score. Lower perplexity indicates smoother overall contextual semantics. The calculation formula for semantic fluency is defined as follows:
$PPL_{f}=sum\left(PPL_{q}\left([c_{l},q,c_{r}]\mid\cdot\right)+PPL_{l}\left([c_{l},q,c_{r}]\mid\cdot\right)\right)$ (6)
where $\text{PPL}_{q}\left([c_{l};q;c_{r}]\mid\cdot\right)$, $\text{PPL}_{l}\left([c_{l};q;c_{r}]\mid\cdot\right)$ represents the perplexity of the whole context for Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023) respectively.
@@
Similarly, for the convenience of calculation, the final score for semantic fluency is designed as follows:
$S_{f}=\left(1-\frac{PPL_{f}}{50}\right)*100\%$ (7)
@@
Quotation Novelty. Integrating novel quotations into established ideas enhances originality and personalizes the expression within academic discourse. To evaluate the extent to which the quote introduces new ideas or unique perspectives to the original context, we also use Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023) to calculate the quotation ppl. Additionally, we utilize the Bing [^3] https://www.bing.com/search engine to determine the number of Search Frequency corresponding to each quotation, applying a log10 transformation to quantify quotation popularity. The calculation formula for quotation popularity is defined as follows:
$Novelty=\frac{PPL(q\mid\cdot)\times 5}{log_{10}(SearchFrequency)}$ (8)
where $PPL(q\mid\cdot)$ refers to the sum ppl for calculating the cited quotation of Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023), and Search Frequency indicates the number of search results obtained by searching the quotation in the Bing search engine.
In the same way, we set a benchmark Novelty value (Novelty=20), which can be considered to have considerable novelty. So the final score for quotation novelty is as follows:
$S_{n}=\left(\frac{Novelty}{20}\right)*100\%$ (9)
@@ 5.1 Dataset Construction
Table 1: The statistics of our knowledge base. For each category, the AvgLen, AvgSearchFreq and AvgNovelty denote the average of the length, the frequency of Bing Search engine and the value of Quotation Novelty respectively.
@@ 6 Quotation-specific Reranking Metric
Quotation matching emphasizes the completion of the quotation itself and its alignment with the subsequent text. This is accomplished by calculating the perplexity (PPL) of the remaining portion of the quotation, given the preceding text and the initial k characters of the quotation. Generally, lower PPL values suggest that the model produces more accurate and coherent quotations. The specific calculation formula is as follows:
@@
where n represents the length of the quote, $q_{t}$ represents the first t characters of the quote, $q_{n-t}$ represents the remaining n-t characters of the quote, and $PPL_{m}$ represent the sum perplexity of Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023) respectively.
@@
Semantic Matching
Semantic matching is concerned with ensuring semantic consistency and logical coherence within the context. This is achieved by predicting the PPL of the subsequent text, given the preceding text and the entire quote. Typically, a lower PPL value indicates that the context with the quote generated by the model is more logically consistent. The specific calculation formula is as Equation (10).
@@ 7 Experiments
ChatGLM3-6B 0.56 0.20 0.73 0.73 0.74 0.59
Qwen1.5-7B-Chat 0.63 0.15 0.71 0.72 0.70 0.58
Qwen1.5-14B-Chat 0.66 0.16 0.69 0.70 0.74 0.60
Qwen1.5-72B-Chat 0.72 0.16 0.70 0.71 0.71 0.60
@@
Mixture-7B-v0.2 0.77 0.08 0.69 0.74 0.58 0.57
Llama2-7B-Chat-hf 0.46 0.15 0.73 0.72 0.77 0.57
Llama2-13B-Chat-hf 0.48 0.15 0.73 0.73 0.78 0.58
Llama2-70B-Chat-hf 0.60 0.11 0.69 0.71 0.66 0.55
@@ 7.1 Experiment Setup
ChatGLM3-6B 0.56 0.20 0.73 0.73 0.74 0.59 0.59 0.13 0.71 0.72 0.71 0.57 0.62 0.13 0.71 0.72 0.72 0.58 0.64 0.16 0.70 0.72 0.71 0.59
Qwen1.5-7B-Chat 0.63 0.15 0.71 0.72 0.70 0.58 0.66 0.13 0.71 0.71 0.75 0.59 0.67 0.13 0.70 0.72 0.73 0.59 0.67 0.13 0.71 0.72 0.73 0.60
Qwen1.5-14B-Chat 0.66 0.16 0.69 0.70 0.74 0.60 0.68 0.17 0.71 0.72 0.71 0.60 0.74 0.18 0.70 0.72 0.69 0.61 0.69 0.18 0.70 0.73 0.72 0.61
Qwen1.5-72B-Chat 0.72 0.16 0.70 0.71 0.71 0.60 0.67 0.21 0.71 0.72 0.70 0.60 0.63 0.18 0.71 0.71 0.70 0.59 0.78 0.20 0.70 0.71 0.71 0.62
@@
Mixture-7B-v0.2 0.77 0.08 0.69 0.74 0.58 0.57 0.82 0.17 0.69 0.75 0.55 0.59 0.82 0.15 0.69 0.74 0.50 0.58 0.77 0.09 0.70 0.73 0.61 0.58
Llama2-7B-Chat-hf 0.46 0.15 0.73 0.72 0.77 0.57 0.46 0.09 0.74 0.72 0.70 0.54 0.44 0.12 0.74 0.73 0.71 0.55 0.49 0.14 0.74 0.73 0.74 0.56
Llama2-13B-Chat-hf 0.48 0.15 0.73 0.73 0.78 0.58 0.44 0.10 0.75 0.72 0.78 0.56 0.50 0.13 0.73 0.70 0.78 0.57 0.45 0.10 0.73 0.70 0.77 0.55
Llama2-70B-Chat-hf 0.60 0.11 0.69 0.71 0.66 0.55 0.65 0.20 0.70 0.71 0.71 0.59 0.70 0.20 0.70 0.72 0.67 0.59 0.75 0.13 0.71 0.71 0.70 0.60
@@
GPT-3.5-turbo 0.62 0.11 0.70 0.73 0.65 0.56 0.72 0.16 0.70 0.74 0.62 0.59 0.73 0.14 0.70 0.74 0.61 0.58 0.76 0.10 0.70 0.72 0.60 0.57
GPT-4o 0.79 0.23 0.71 0.74 0.62 0.62 0.75 0.24 0.70 0.74 0.63 0.61 0.80 0.23 0.70 0.74 0.60 0.62 0.83 0.22 0.72 0.74 0.64 0.63
Table 5: Comparison of performance of various models on our evaluation system for QG tasks in in Naive-0-shot, Naive-1-shot, Naive-2-shot and Naive-cot settings. In these naive experimental setup, our experiment does not employ RAG or rerank metrics. Instead, it relies solely on a specifically designed prompt to enable the models to execute the QG task. Overall, the GPT-4 perform better than other models. The prompt for each setting is detailed in the Appendix.
@@ 7.2 Results
Overall Performance
We conduct experiments on models of different ranges and sizes on our benchmark, and the results are shown in Table 3. Among the evaluated models, GPT-4 outperforms the other models, followed by Qwen1.5-72B-Chat. Despite varying parameter sizes (ranging from 6B to 72B), all models demonstrate suboptimal performance on the quotation generation task. Even the best-performing model, GPT-4, achieves an average score of no more than 0.62 across our five evaluation metrics, highlighting the critical need to address the quotation hallucination problem. Notably, our Quotation-specific Reranking method achieves the best results in each indicator, demonstrating the concise effectiveness of our proposed reranking metric.
Comparison between Model Types
Furthermore, the performance comparison between the Chinese-oriented group and the English-oriented group on the Chinese-English benchmark reveals no significant differences, suggesting that the model’s quotation ability is not language-dependent. Overall, the current opensource small to large-scale models exhibit a relatively small performance gap compared to close-source large-scale models, indicating the universality of the issue of quotation hallucination in LLMs.
@@
In addition, we conduct further analysis on different model sizes. After applying our method, both Authenticity and Credibility should ideally reach 1, which highlights the method’s effectiveness in assisting models with limited cited knowledge, thereby mitigating quotation hallucination. Moreover, within the same series, larger models tend to show improved performance. This indicates that larger models have richer quotation memory and stronger instruction-following capabilities.

### v2

@@ 1 Introduction
Figure 1: An example of prevalent issues in Quotation Generation (QR) by LLMs. In QR tasks, LLMs often fabricate sentences, leading to quotation hallucination. Additionally, the generated quotes may not align with the context, resulting in contextual inconsistency and semantic incoherence. Finally, the sentences produced by LLMs tend to be overly common, resulting in a lack of novelty in quotations.
@@ 4 Evaluation System for QG
Quotation Authenticity. Authenticity of quotations is crucial as it ensures the reliability and credibility of information Kington et al. (2021). To verify the authenticity of the quoted celebrity quotes, we first search the quotation database for the information corresponding to the quote. If the database contains the information, we use the corresponding information to make a judgment. If not, we use different search engines (such as Google Scholar [^1] https://scholar.google.com/ and Baidu Scholar [^2] https://xueshu.baidu.com/) to recall the corresponding search results. Previous studies Han et al. (2024) have shown that GPT-4o OpenAI (2022) has excellent simple extraction capabilities, and the extraction task based on this study only has two fields, author and source. Therefore, we use GPT-4o to extract the corresponding field information, and then compare the results of different search engines. If the field information is different, manual comparison is required. For extraction details and validity of GPT-4o, please refer to Appendix C. Finally, based on the extracted information, we verify whether the quote genuinely originates from the specific celebrity. The final score is defined as follows:
@@
Quotation Credibility. Generally speaking, in the context of quoting, the source of the quote will be mentioned, such as the author, classic literature, or other sources. Ensuring consistency between the citation and the mentioned author or source is crucial for maintaining contextual coherence and information accuracy. In order to confirm whether the citation meets the source restriction mentioned in the context, our study first extracts the source restriction of the context, and then compares and analyzes it with the extraction result of the previous indicator. If the source matches, the citation is marked as trustworthy, as shown in Fig.1. The final score for quotation credibility is defined as follows:
@@
Semantic Matching. Improper quotation may lead to misunderstandings or misinterpretations of the original meaning, thereby weakening the effectiveness and persuasiveness of the argument Quora (2020). Perplexity is a common metric in NLP, used to assess a language model’s predictive capability for text. Hence, we calculate the PPL of subsequent text based on a given prior text and quotation to evaluate the consistency between the quotation and its context. If the evaluation score is low, it implies that the citation aligns well with the following context in terms of semantics; otherwise, the rationality of the citation should be reconsidered. The formula is as follows:
$PPL_{m}=PPL\left(c_{r}\mid[c_{l};q]\right)$ (4)
where $c_{l}$ stands for the previous text, $q$ stands for the quotation, and $c_{r}$ stands for the following text.
To simplify computation, we normalize the PPL values to a range between 0 and 1. Given that a lower PPL indicates a higher degree of semantic alignment, we utilize an inverted Sigmoid function. The final calculation formula is as follows:
$S_{m}=\dfrac{1}{1+e^{k_{m}(PPL_{m}-\mu_{m})}}*100\%$ (5)
where $\mu_{m}$ represents the mean value of $PPL_{m}$, which is 35.243, and $k_{m}$ is determined using an empirical formula, yielding a value of 0.053. See the Appendix D for the specific calculation details.
@@
Semantic Fluency. After quotation, it is necessary to ensure that the entire context is fluent and coherent to maintain semantic consistency and logical integrity Krumm et al. (2020). This study calculates the PPL of the entire context to measure the textual fluency of the overall context after inserting quotations. Lower perplexity indicates smoother overall contextual semantics. The calculation formula for semantic fluency is as follows:
$PPL_{f}=PPL_{q}\left([c_{l},q,c_{r}]\mid\cdot\right)$ (6)
where $c_{l}$ stands for the previous text, $q$ stands for the quotation, and $c_{r}$ stands for the following text.
@@
Similarly, for normalizing the PPL values into a range from 0 to 1, the final score for semantic fluency is designed as follows:
$S_{f}=\dfrac{1}{1+e^{k_{f}(PPL_{f}-\mu_{f})}}*100\%$ (7)
where $\mu_{f}$ represents the mean value of $PPL_{f}$, which is 16.470, and $k_{f}$ is determined using an empirical formula, yielding a value of 0.500.
@@
Quotation Novelty. Integrating novel quotations into established ideas enhances originality and personalizes the expression within academic discourse Savov (2021). To evaluate the extent to which the quote introduces new ideas or unique perspectives to the original context, we utilize the Bing [^3] https://www.bing.com/search engine to determine the number of Search Frequency corresponding to each quotation, applying a log10 transformation to quantify quotation popularity. In addition, to mitigate potential biases in search results, we also incorporate the quoted PPL value for supplementation. As a lower PPL indicates a higher frequency of text occurrence, it is inversely correlated with search frequency. Therefore, the formula is as follows:
$\text{Novelty}=\frac{PPL(q\mid\cdot)}{log_{10}(\text{Search Frequency})}$ (8)
where Search Frequency indicates the number of search results obtained by searching the quotation in the Bing search engine. In order to map the PPL value to a range of 0 to 1, since higher novelty means higher score, the positive sigmoid function is adopted here, and the final score is as follows:
$S_{n}=\dfrac{1}{1+e^{-k_{n}(Novelty-\mu_{n})}}*100\%$ (9)
where $\mu_{n}$ represents the mean value of Novelty, which is 10.67, and $k_{n}$ is determined using an empirical formula, yielding a value of 0.253.
@@ 5.1 Dataset Construction
Table 1: The statistics of our knowledge base. For each category, the AvgLen, AvgSearchFreq and AvgNovelty denote the average of the length, the frequency of Bing Search engine, and the value of Quotation Novelty respectively.
@@ 5.2 Dataset Statistics
5.3 Quality Assessment by Human
After constructing the dataset, we manually check its quality. For each component, we randomly select 100 quotes and engage three annotators to verify their validity. The annotators use search engines  [^10] https://www.bing.com/ to locate references and evaluate both the authenticity of the quotes and the accuracy of their attributed authors and sources. Only quotes that satisfy both criteria are included in the final dataset. The final results are determined through a majority voting process. In the English, Standard Chinese and Classical Chinese components, 99, 97 and 98 quotations respectively met the established criteria. These results confirm the high quality of the dataset, which is derived from trustworthy sources such as published books and reputable citation websites.
@@ 6 Quotation-specific Reranking Metric
Quotation matching emphasizes the completion of the quotation itself and its alignment with the subsequent text MacLaughlin and Smith (2021). This is accomplished by calculating the PPL of the remaining portion of the quotation, given the preceding text and the initial k characters of the quotation. Generally, lower PPL values suggest that the model produces more accurate and coherent quotations. The specific calculation formula is as follows:
@@
where $n$ represents the length of the quote, $q_{t}$ represents the first $t$ characters of the quote, $q_{n-t}$ represents the remaining $n-t$ characters of the quote.
Semantic Matching
Semantic matching is concerned with ensuring semantic consistency and logical coherence within the context. This is achieved by predicting the PPL of the subsequent text, given the preceding text and the entire quote. A lower PPL value means that the quotation is more semantically consistent with the following context. The calculation formula is as Equation (4).
@@ 7 Experiments
ChatGLM3-6B 0.56 0.20 0.72 0.73 0.71 0.58
Qwen1.5-7B-Chat 0.63 0.15 0.72 0.68 0.71 0.58
Qwen1.5-14B-Chat 0.66 0.16 0.72 0.69 0.73 0.60
Qwen1.5-72B-Chat 0.72 0.16 0.71 0.71 0.67 0.60
Deepseek-R1 0.70 0.39 0.72 0.76 0.49 0.61
@@
Mixture-7B-v0.2 0.77 0.08 0.70 0.74 0.55 0.57
Llama2-7B-Chat-hf 0.46 0.15 0.73 0.73 0.74 0.56
Llama2-13B-Chat-hf 0.48 0.15 0.74 0.72 0.74 0.56
Llama2-70B-Chat-hf 0.60 0.11 0.69 0.67 0.62 0.55
@@ 7.1 Experiment Setup
Models for PPL Calculation
We employ two advanced language models, Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023). These models are used to compute the PPL of the context given the preceding text. Subsequently, the average PPL values calculated by the two models are taken as final PPL values, which balances the judgments of the two models and reduces potential bias introduced by any single model. Since larger models tend to produce lower PPL for the same text, we recommend using the same PPL calculation models in this study when evaluating QG tasks.
@@
ChatGLM3-6B 0.56 0.20 0.72 0.73 0.71 0.58 0.59 0.13 0.72 0.68 0.67 0.56 0.62 0.13 0.72 0.68 0.68 0.57 0.64 0.16 0.71 0.69 0.67 0.57
Qwen1.5-7B-Chat 0.63 0.15 0.72 0.68 0.71 0.58 0.66 0.13 0.72 0.70 0.71 0.59 0.67 0.13 0.71 0.70 0.69 0.58 0.67 0.13 0.72 0.69 0.69 0.59
Qwen1.5-14B-Chat 0.66 0.16 0.72 0.69 0.73 0.60 0.68 0.17 0.72 0.67 0.71 0.60 0.74 0.18 0.71 0.71 0.65 0.60 0.69 0.18 0.72 0.73 0.68 0.60
Qwen1.5-72B-Chat 0.72 0.16 0.71 0.71 0.67 0.60 0.67 0.21 0.72 0.72 0.67 0.60 0.63 0.18 0.72 0.71 0.65 0.58 0.78 0.20 0.70 0.71 0.65 0.61
Deepseek-R1 0.70 0.39 0.72 0.76 0.49 0.61 0.67 0.38 0.71 0.71 0.54 0.60 0.71 0.38 0.71 0.72 0.54 0.62 0.77 0.35 0.71 0.74 0.54 0.62
@@
Mixture-7B-v0.2 0.77 0.08 0.70 0.74 0.55 0.57 0.82 0.17 0.71 0.75 0.52 0.59 0.82 0.15 0.70 0.75 0.46 0.58 0.77 0.09 0.71 0.73 0.58 0.58
Llama2-7B-Chat-hf 0.46 0.15 0.73 0.73 0.74 0.56 0.46 0.09 0.73 0.71 0.66 0.53 0.44 0.12 0.73 0.74 0.67 0.54 0.49 0.14 0.74 0.73 0.70 0.56
Llama2-13B-Chat-hf 0.48 0.15 0.74 0.72 0.74 0.56 0.44 0.10 0.74 0.72 0.74 0.56 0.50 0.13 0.73 0.68 0.74 0.57 0.45 0.10 0.73 0.67 0.74 0.55
Llama2-70B-Chat-hf 0.60 0.11 0.69 0.67 0.62 0.55 0.65 0.20 0.71 0.66 0.67 0.58 0.70 0.20 0.71 0.69 0.63 0.59 0.75 0.13 0.71 0.68 0.66 0.59
@@
GPT-3.5-turbo 0.62 0.11 0.71 0.72 0.62 0.56 0.72 0.16 0.71 0.75 0.59 0.59 0.73 0.14 0.71 0.74 0.57 0.58 0.76 0.10 0.71 0.70 0.58 0.57
GPT-4o 0.79 0.23 0.71 0.74 0.58 0.61 0.75 0.24 0.70 0.74 0.60 0.61 0.80 0.23 0.71 0.76 0.57 0.62 0.83 0.22 0.71 0.73 0.60 0.62
Table 5: Comparison of performance of various models on our evaluation system for QG tasks in in Naive-0-shot, Naive-1-shot, Naive-2-shot and Naive-cot settings. In these naive experimental setup, our experiment does not employ RAG or rerank metrics. Instead, it relies solely on a specifically designed prompt to enable the models to execute the QG task. The prompt for each setting is detailed in the Appendix  LABEL:naive.
@@ 7.2 Results
We conduct experiments on models of different ranges and sizes on our benchmark, and the results are shown in Tab. 3. For more detailed analysis, please refer to Appendix  E.
Severity of Quotaiton Hallucination
The results show that more than half of the citations generated by LLaMA2-13B-Chat are not genuine quotes. Furthermore, despite varying parameter sizes, all models demonstrate suboptimal performance on the QR task, especially on the $S_{c}$ metric. Even the best-performing model, GPT-4o, only scores 0.23 on the $S_{c}$ indicator, highlighting the critical need to address the quotation hallucination problem.
Performance of Quotation-specific Reranking Metric
Notably, our Quotation-specific Reranking method achieves the best results in each indicator, demonstrating the effectiveness of our method. Since our method retrieves the most relevant and appropriate citations from the quotation database, it ensures the authenticity and credibility of the citations. Therefore, both $S_{a}$ and $S_{c}$ are equal to 1. In addition, our method can effectively improve the novelty of citations and alleviate the problem of generating common citations with LLMs.
@@
We conduct further analysis on different model sizes. Within the same series, larger models tend to show improved performance. This indicates that larger models have richer quotation memory and stronger instruction-following capabilities.